# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ Appendix I Complete Ideal Functionality for Profile P2
Since $\mathsf{result}$ is derivable from $\sigma$ (a valid signature implies $\mathsf{result}=1$), no additional information is leaked.
@@
Table 14: Detailed round and communication breakdown for Profile P2 r0-check (original 8-round version).
@@ 5 Output
The combiner’s role is purely message routing with simultaneous release. We show the adversary’s view is identical:
@@
  • In $\mathsf{H}_{0}$: Combiner collects all $(Com,value,rand)$ tuples, releases simultaneously
  • In $\mathsf{H}_{0.5}$: $\mathcal{F}_{\mathsf{Combiner}}$ does the same
  • Adversary sees: all commitments (at send time), all openings (at release time)
  • Timing of release is identical (after all collected)
@@
  • Adversary controls combiner directly
  • Receives honest parties’ messages as in real protocol
  • Any deviation (selective release, modification) is detectable:
@@
      Commitment mismatch: immediate abort
@@
      Value modification: SPDZ MAC check fails
  • Functionality captures this via abort mechanism
Since the adversary’s view and available actions are identical in both hybrids:
@@
$\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}+\epsilon_{\mathsf{combiner}}=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
@@
$\epsilon=0+\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}+0+0+0+0<2^{-11}$
@@
  • Rounds 6-7 $\to$ Round 4: Parallel comparison + AND tree in single round
  • Round 8 $\to$ Round 5: Output (unchanged)
@@
The adversary $\mathcal{A}$ can:
@@
  • Control message delivery within each phase
@@
Functionality P.3 ($\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ – Semi-Async r0-Check).
Extends $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ (Definition 6.4) with pre-signing and timing.
@@
Protocol P.4 ($\Pi_{r_{0}}^{\mathsf{P3+}}$ – P3+ Semi-Async Protocol).
@@
      Sample $\mathsf{set\_id}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
@@
      Compute commitment $\mathsf{Com}_{i}=H(\mathbf{y}_{i}\|\mathbf{w}_{i}\|\mathsf{set\_id})$
@@
    Send $(\mathsf{set\_id},\mathbf{w}_{i},\mathsf{Com}_{i})$ to combiner
@@
    If $\mathsf{result}=1$: The 2PC circuit also outputs the hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$ (computed inside the circuit using $c\mathbf{s}_{2}=\mathbf{S}_{1}+\mathbf{S}_{2}$); combiner outputs signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
@@
Simulator P.5 ($\mathcal{S}^{\mathsf{P3+}}$ for $\Pi_{r_{0}}^{\mathsf{P3+}}$).
@@
      Sample random $\tilde{\mathbf{w}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{k\times n}$
@@
      Sample random $\tilde{\mathsf{Com}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
@@
    Justification: Commitment hides $\mathbf{y}_{i}$ (preimage resistance). Simulated $\mathbf{w}_{i}$ is uniform, same as real (since $\mathbf{y}_{i}$ is uniform and $\mathbf{A}$ is public).
@@
      Sample random $\tilde{\mathbf{z}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{l\times n}$
@@
      Sample random $\tilde{\mathbf{V}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{k\times n}$
@@
      Send $(\tilde{\mathbf{z}}_{i},\tilde{\mathbf{V}}_{i})$ to adversary (via combiner)
@@
    Justification: By Theorem 4.4 (statistical privacy), the Shamir nonce DKG ensures each honest party’s response $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ has high conditional min-entropy ($\geq 5\times$ secret key entropy for $|S|\leq 17$) from the adversary’s perspective without computational assumptions. The $\mathbf{V}_{i}$ values are uniform by PRF-based mask hiding (Lemma 4.9). Simulated (uniform) values have negligible statistical distance from real values.
@@
    $\mathcal{S}$ extracts $(\mathbf{w},\mathbf{S}_{1})$ from adversary’s garbled circuit (extractability)
@@
      Generate simulated garbled circuit that evaluates to $\mathsf{result}$
@@
      This is possible by simulatability of garbled circuits [BHR12]
  6. 6.
    Justification: Garbled circuit simulatability: given output $f(x)$, can simulate garbled circuit indistinguishable from real.
@@
  • $\mathsf{H}_{3}$: Simulate pre-signing as random
  • $\mathsf{H}_{4}$: Simulate responses as random
@@
Lemma P.7 (Pre-signing Simulation).
Proof P.8.
Pre-signing values visible to adversary:
  • $\mathbf{w}_{i}=\mathbf{A}\cdot\mathbf{y}_{i}$: Since $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{l\times n}$ uniform and $\mathbf{A}$ is full rank (with overwhelming probability), $\mathbf{w}_{i}$ is uniform over $\mathbb{Z}_{q}^{k\times n}$.
  • $\mathsf{Com}_{i}=H(\mathbf{y}_{i}\|\mathbf{w}_{i}\|\mathsf{set\_id})$: By random oracle model, $\mathsf{Com}_{i}$ is uniform $\{0,1\}^{256}$.
Simulated values $(\tilde{\mathbf{w}}_{i},\tilde{\mathsf{Com}}_{i})$ are sampled uniformly from same distributions. Thus:
@@
$|\Pr[\mathsf{H}_{3}=1]-\Pr[\mathsf{H}_{4}=1]|\leq\mathsf{negl}(\kappa)$ (statistical indistinguishability)
Proof P.10.
For honest signer $P_{i}$:
@@
By Theorem 4.4 (statistical privacy), the Shamir nonce DKG ensures that the adversary observes only T1T-1 evaluations of each honest party’s degree-(T1)(T\!-\!1) nonce polynomial, leaving one degree of freedom that provides high conditional min-entropy. Due to the bounded-nonce constraint, $\mathbf{z}_{i}$ is not perfectly $R_{q}^{\ell}$-uniform but has min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits, vastly exceeding the secret key entropy. The $\mathbf{V}_{i}$ values contain pairwise masks $\mathbf{m}_{i}^{(s2)}$ derived from PRF seeds shared only with other honest signers; against an adversary controlling fewer than all honest signers, at least one honest signer contributes an unknown mask component that acts as a one-time pad.
Simulated $\tilde{\mathbf{z}}_{i},\tilde{\mathbf{V}}_{i}$ are sampled uniformly, matching the real distribution.
@@
Protocol $\Pi_{r_{0}}^{\mathsf{P3+}}$ UC-realizes $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ in the $(\mathcal{F}_{\mathsf{OT}},\mathcal{F}_{\mathsf{GC}})$-hybrid model against static malicious adversaries corrupting up to T1T-1 signers and at most one of $\{\mathsf{CP}_{1},\mathsf{CP}_{2}\}$, with security-with-abort.
@@
$|\Pr[\mathsf{REAL}=1]-\Pr[\mathsf{IDEAL}=1]|\leq\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}<2^{-\kappa}$
where $\kappa$ is the computational security parameter.
Proof P.12.
@@
$|\Pr[\mathsf{H}_{0}]-\Pr[\mathsf{H}_{5}]|$ $\leq\sum_{i=0}^{4}|\Pr[\mathsf{H}_{i}]-\Pr[\mathsf{H}_{i+1}]|$
$=\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}+0+0+0$

### v5

@@ Appendix I Complete Ideal Functionality for Profile P2
For successful attempts, $\mathsf{result}=1$ is implied by the existence of $\sigma$ and leaks no new information. For failed attempts ($\mathsf{result}=0$, no $\sigma$ produced), the result bit is determined by the public broadcast $(\tilde{c},\mathbf{w})$ and the corrupted parties’ $r_{0}$-check shares $\{c\mathbf{s}_{2,j}\}_{j\in C}$: the simulator computes
$\mathbf{r}_{0}^{(C)}=\mathsf{HighBits}\!\left(\sum_{j\in C}\lambda_{j}\,c\,\mathbf{s}_{2,j}+\text{honest aggregate contribution}\right)$
and determines the result by whether $\|\mathbf{r}_{0}\|_{\infty}<\gamma_{2}-\beta$. Since the honest contribution is masked (mask-hiding, Lemma 4.15), the simulator learns only the aggregate result from $\mathcal{F}_{r_{0}\text{-check}}$, not the individual honest shares. The failed-attempt result bits can thus be provided to the simulator by $\mathcal{F}_{r_{0}\text{-check}}$ without leaking additional information beyond what is already in $\mathcal{Z}$’s view.
@@
Remark M.5 (Nonce DKG Simulation for P2).
The P2 UC proof above is scoped to the r0-check subprotocol $\Pi_{r_{0}}$ realizing $\mathcal{F}_{r_{0}\text{-check}}$. The nonce DKG simulation for the full P2 protocol follows the same structure as P1/P3+: the simulator generates honest nonce shares from simulated degree-(T1)(T-1) polynomials, producing evaluations with identical distribution to the real protocol. Specifically, $\mathcal{S}$ runs the nonce DKG on behalf of honest parties (sampling the high-degree coefficients uniformly from $R_{q}$), provides consistent evaluations to corrupted parties, and feeds the resulting nonce shares into the r0-check simulation above. By the UC composition theorem, the full P2 signing protocol (nonce DKG + response computation + r0-check) securely composes these subprotocols.
@@
Table 14: Round and communication breakdown for Profile P2 r0-check (8-round version). The AND tree requires m1m-1 Beaver-triple openings, giving $O(m)$ bits; the $O(\log m)$ figure denotes circuit depth, not communication.
@@
Timeout (Corrupted Combiner):
Upon sending $(\mathsf{ALL\_COLLECTED},\mathsf{sid})$ to the combiner, the functionality starts a timer of $\Delta_{\mathsf{timeout}}$ rounds (a global synchrony parameter). If $(\mathsf{RELEASE},\mathsf{sid})$ is not received within $\Delta_{\mathsf{timeout}}$ rounds:
  1. 1.
    Send $(\mathsf{ABORT},\mathsf{sid})$ to all parties $P_{i}$
  2. 2.
    Discard all collected state
Remark: This timeout prevents a corrupted combiner from indefinitely delaying the release. In Profiles P1 and P3+, the combiner role is enforced by a TEE or 2PC, so the timeout is triggered only if the trusted component is unavailable (which is detected). In Profile P2, all parties are online during the signing window, so $\Delta_{\mathsf{timeout}}$ is a protocol parameter (e.g., 30 seconds).
@@ 5 Output
The combiner performs no cryptographic operations: it collects incoming messages and forwards them once all have arrived (or times out). In particular, it does not compute any MACs, encryptions, or protocol-specific transformations on the data. Its behavior is a deterministic function of its input messages and the arrival timing. We show the adversary’s view is bytewise identical in $\mathsf{H}_{0}$ and $\mathsf{H}_{0.5}$:
@@
  • In $\mathsf{H}_{0}$: The combiner collects $(\mathsf{Com}_{i},\mathsf{value}_{i},\mathsf{rand}_{i})$ tuples from all $N$ parties and releases them simultaneously only after all are received.
  • In $\mathsf{H}_{0.5}$: $\mathcal{F}_{\mathsf{Combiner}}$ executes the same algorithm: collect-then-release.
  • The adversary’s incoming messages (the individual party tuples) are produced by the real parties’ code in $\mathsf{H}_{0}$ and by the simulator (which replicates the same code) in $\mathsf{H}_{0.5}$—producing identical bit-strings.
  • The adversary’s outgoing view (released tuples) is the same set of collected messages, released at the same logical time.
  • Since the combiner is honest, the adversary’s view of the combiner’s internal state is empty (it receives only the released outputs). The released outputs are identical. Hence the adversary’s view is bytewise equal in both hybrids.
@@
  • The adversary controls the combiner directly and receives honest parties’ messages in both hybrids via identical delivery.
  • Any deviation by the combiner (selective release, modification) triggers a detectable abort:
@@
      Value modification: SPDZ MAC check at the recipient parties fails with probability 11/q1-1/q per coefficient; honest parties abort and the functionality’s abort interface captures this.
@@
      Selective withholding: honest parties’ timeout logic triggers an abort after $\mathsf{deadline}$, identical in both $\mathsf{H}_{0}$ and $\mathsf{H}_{0.5}$.
  • In $\mathsf{H}_{0.5}$, $\mathcal{F}_{\mathsf{Combiner}}$ replicates the same abort interface. Since the honest parties’ messages delivered to the corrupted combiner are generated by the same real-protocol code in both hybrids (the combiner’s corruption does not affect honest-party message generation), the adversary’s view is again bytewise identical.
Since the combiner performs no cryptographic operations and its behavior is a deterministic routing function, the adversary’s view in $\mathsf{H}_{0}$ is perfectly equal to its view in $\mathsf{H}_{0.5}$ (not merely computationally indistinguishable). Thus:
@@
$\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}+\epsilon_{\mathsf{combiner}}=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-53}$
(with 64-bit MAC field; minimal configuration gives $\epsilon<2^{-11}$).
@@
$\epsilon=0+\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}+0+0+0+0<2^{-53}$
@@
  • Rounds 6-7-8 $\to$ Rounds 4-5: Parallel comparison circuits and AND aggregation execute simultaneously in Round 4 (saving 1 round); output is broadcast in Round 5 (total: 3 original rounds $\to$ 2 new rounds)
@@
We adopt the synchronous UC model [Katz07] where all parties share a global clock. The adversary $\mathcal{A}$ can:
@@
  • Control message scheduling within each phase, but cannot advance the clock past the signer response deadline $\Delta$
@@
Theorem P.12 applies to any protocol meeting this interface.
@@
Definition P.3 (Ideal Functionality $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ (UC Proof Compact Form)).
The functionality $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ captures a two-party computation of the r0-check predicate. It receives private input $\mathbf{S}_{j}\in\mathbb{Z}_{q}^{m}$ from each computation party $\mathsf{CP}_{j}$ ($j\in\{1,2\}$), reconstructs $c\mathbf{s}_{2}=\mathbf{S}_{1}+\mathbf{S}_{2}\bmod q$ (private inputs satisfy this additive split; in the concrete protocol, $\mathbf{S}_{j}=\sum_{i\text{ assigned to }\mathsf{CP}_{j}}\lambda_{i}\cdot c\mathbf{s}_{2,i}$), computes $\mathbf{w}^{\prime}=\mathbf{w}-c\mathbf{s}_{2}$ (using the public $\mathbf{w}$ agreed upon during challenge derivation), evaluates $\mathsf{result}=\bigwedge_{j}(|\mathsf{LowBits}(\mathbf{w}^{\prime}_{j},2\gamma_{2})|_{\infty}<\gamma_{2}-\beta)$, and outputs only $\mathsf{result}$ (1 bit) and the hint $\mathbf{h}=\mathsf{MakeHint}(\mathbf{w}^{\prime},\mathbf{w},2\gamma_{2})$ to the combiner. Neither $c\mathbf{s}_{2}$ nor $\mathbf{w}^{\prime}$ is revealed outside $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$.
Functionality P.4 ($\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ – Semi-Async r0-Check).
Extends $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ (Definition P.3) with pre-signing and timing.
@@
Protocol P.5 ($\Pi_{r_{0}}^{\mathsf{P3+}}$ – P3+ Semi-Async Protocol).
@@
      Sample $\mathsf{set\_id}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ (pre-signing set identifier)
@@
      Sample commitment randomness $r_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
    5. (e)
      Compute commitment $\mathsf{Com}_{i}=H(\texttt{"com"}\|\mathbf{y}_{i}\|\mathbf{w}_{i}\|r_{i})$
@@
    Send $(\mathsf{set\_id},\mathbf{w}_{i},\mathsf{Com}_{i})$ to combiner $\triangleright$ $r_{i}$ is not sent; kept secret by signer
@@
    If $\mathsf{result}=1$: The 2PC circuit outputs the hint
    $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\;\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},\;2\gamma_{2})$
    (computed inside the circuit using $c\mathbf{s}_{2}=\mathbf{S}_{1}+\mathbf{S}_{2}$); the combiner outputs signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
@@
Simulator P.6 ($\mathcal{S}^{\mathsf{P3+}}$ for $\Pi_{r_{0}}^{\mathsf{P3+}}$).
@@
Simulating Key DKG:
  1. 1.
    $\mathcal{S}$ receives the corrupted parties’ key shares $\{(\mathbf{s}_{1,j},\mathbf{s}_{2,j})\}_{j\in C}$ from the adversary (via corruption; in P3+ the key DKG is modeled as a trusted-dealer setup phase, so the simulator acts as the trusted dealer and generates all shares directly).
  2. 2.
    $\mathcal{S}$ runs the ideal key-generation functionality on behalf of all honest parties: $\mathcal{S}$ samples a fresh degree-(T1)(T-1) Shamir polynomial for each component of $(\mathbf{s}_{1},\mathbf{s}_{2})$, consistent with the corrupted-party shares received in step 1, and computes the honest-party evaluations $\{(\mathbf{s}_{1,i},\mathbf{s}_{2,i})\}_{i\notin C}$ directly from these polynomials.
  3. 3.
    $\mathcal{S}$ stores all shares $\{(\mathbf{s}_{1,i},\mathbf{s}_{2,i})\}_{i=1}^{N}$ for use in the response simulation.
  4. 4.
    Justification: The simulator acts as the DKG simulator, generating the full Shamir polynomials during key setup. It does not infer honest shares post-hoc from $\mathbf{t}_{1}$ (which is a lossy commitment and does not uniquely determine $(\mathbf{s}_{1},\mathbf{s}_{2})$). Rather, the polynomials are fixed at key-generation time by $\mathcal{S}$’s own choices, and all evaluations $\{(\mathbf{s}_{1,i},\mathbf{s}_{2,i})\}_{i=1}^{N}$ are therefore directly available to $\mathcal{S}$. When |C|<T1|C|<T-1, there are multiple polynomials consistent with the corrupted shares; $\mathcal{S}$ uses the one it sampled.
@@
      Sample $\tilde{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{R}_{q}^{\ell}$ (simulated nonce share)
@@
      Compute $\tilde{\mathbf{w}}_{i}=\mathbf{A}\cdot\tilde{\mathbf{y}}_{i}$ (same computation as real protocol)
@@
      Sample $\tilde{r}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ (commitment randomness); program the random oracle so that $H(\mathsf{``com^{\prime\prime}}\|\tilde{\mathbf{y}}_{i}\|\tilde{\mathbf{w}}_{i}\|\tilde{r}_{i}):=\tilde{\mathsf{Com}}_{i}$ where $\tilde{\mathsf{Com}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$; store $(\tilde{\mathbf{y}}_{i},\tilde{\mathbf{w}}_{i},\tilde{r}_{i})$ for potential blame-protocol opening
    4. (d)
@@
    Justification: Commitment hides $\mathbf{y}_{i}$ (random oracle). Simulated $\tilde{\mathbf{w}}_{i}=\mathbf{A}\cdot\tilde{\mathbf{y}}_{i}$ has the same distribution as the real $\mathbf{w}_{i}=\mathbf{A}\cdot\mathbf{y}_{i}$: the real $\mathbf{y}_{i}=F(i)$ for $i\neq 0$ is $\mathcal{R}_{q}^{\ell}$-uniform because the degree-11-through-(T1)(T\!-\!1) coefficients of $F$ are sampled uniformly (Theorem Q.3), making each non-zero evaluation uniform by linearity; thus $\tilde{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{R}_{q}^{\ell}$ matches this distribution exactly (SD =0=0). The stored $\tilde{r}_{i}$ enables blame-protocol commitment opening.
@@
      Compute $\mathbf{z}_{i}=\tilde{\mathbf{y}}_{i}+c\cdot\mathbf{s}_{1,i}$ using the nonce shares $\tilde{\mathbf{y}}_{i}$ from the simulated pre-signing phase and key shares $\mathbf{s}_{1,i}$ from the simulated key DKG
@@
      Compute $\mathbf{V}_{i}=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$ using the simulated key shares and PRF masks
@@
      Send $(\mathbf{z}_{i},\mathbf{V}_{i})$ to adversary (via combiner)
@@
    Justification: The simulator computes responses using the same deterministic function as the real protocol, applied to the simulated DKG values. Since the simulated DKG produces values with identical distribution to the real DKG (Lemma P.8), the resulting $\mathbf{z}_{i}=\tilde{\mathbf{y}}_{i}+c\cdot\mathbf{s}_{1,i}$ is perfectly indistinguishable from the real value (SD =0=0, since $\tilde{\mathbf{y}}_{i}$ and $\mathbf{y}_{i}$ have identical distribution). The masked key-share component $\mathbf{V}_{i}=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$ uses PRF-derived masks $\mathbf{m}_{i}^{(s2)}$; the simulator uses the actual pairwise PRF seeds established during key setup simulation, computing the same mask values $\mathbf{m}_{i}^{(s2)}=\mathsf{PRF}(\mathsf{seed}_{h,i},\mathsf{sid})$ as the real protocol. No PRF replacement is needed; the simulation is bit-for-bit identical to the real protocol for this component. The overall gap is therefore 0.
@@
    $\mathcal{S}$ extracts $\mathbf{S}_{1}$ from adversary’s garbled circuit inputs (using garbled-circuit extractability [LP07]); $\mathbf{w}$ is already public (broadcast by the combiner)
@@
      Generate a simulated garbled circuit $\tilde{C}$ together with all input wire labels for $(\mathbf{w},\mathbf{S}_{1})$ (garbler’s inputs) and the wire labels for $\mathbf{S}_{2}$ selected by $\mathsf{CP}_{2}$’s OT queries, consistent with the output $\mathsf{result}$, using the BHR12 simulator for the full two-input garbled circuit [BHR12].
  6. 6.
    Justification: Garbled circuit simulatability (BHR12 privacy definition): given the output $f(\mathbf{w},\mathbf{S}_{1},\mathbf{S}_{2})$ and the evaluator’s input labels (for $\mathbf{S}_{2}$, known from OT), the BHR12 simulator produces a simulated circuit $\tilde{C}$ plus all evaluator-side input labels computationally indistinguishable from the real garbled circuit. The garbler’s (CP1’s) input wire labels for $(\mathbf{w},\mathbf{S}_{1})$ are included in the simulated output; these are internally consistent with $\tilde{C}$ by construction.
Simulating 2PC (Both CPs Honest): When both $\mathsf{CP}_{1}$ and $\mathsf{CP}_{2}$ are honest, the adversary observes only:
  • Pre-signing commitments $\mathsf{Com}_{i}$ (simulated uniformly above, SD =0=0).
  • The broadcast challenge $c$ (determined by the ideal functionality).
  • Responses $(\mathbf{z}_{i},\mathbf{V}_{i})$ for honest $i\notin C$ (simulated above: $\mathbf{z}_{i}$ with SD =0=0; $\mathbf{V}_{i}$ with SD =0=0 since the simulator uses the actual PRF seeds established during key setup, producing bit-for-bit identical mask values).
  • The final result bit (provided by $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$).
The 2PC OT and GC transcripts are exchanged solely between the two honest CPs and are never delivered to the adversary. $\mathcal{S}$ need not simulate these messages; the hybrid sequence $\mathsf{H}_{0}\to\mathsf{H}_{1}$ (replace OT) $\to\mathsf{H}_{2}$ (replace GC) captures the honest-CP case via ideal-functionality substitution.
Simulating Timeout / Abort:
  1. 1.
    If the adversary delays honest signers’ responses past the deadline $\mathsf{deadline}$ (e.g., by withholding the challenge or blocking network messages):
    1. (a)
      $\mathcal{S}$ detects no challenge delivered to honest signers within $\mathsf{deadline}-\Delta$ rounds.
@@
      $\mathcal{S}$ sends $(\mathsf{ABORT},\mathsf{sid})$ to $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$.
    3. (c)
      $\mathcal{S}$ delivers $(\mathsf{abort},\mathsf{sid})$ to all honest parties in the simulated execution.
  2. 2.
    Justification: The functionality enforces “if $\mathsf{now}>\mathsf{deadline}$: reject” (see Response Phase, condition 1). When the adversary induces timeout, both ideal and simulated executions abort identically; the simulated abort message matches the real abort, giving indistinguishability.
@@
  • $\mathsf{H}_{3}$: Simulate pre-signing (sample $\tilde{\mathbf{y}}_{i}$, compute $\tilde{\mathbf{w}}_{i}=\mathbf{A}\tilde{\mathbf{y}}_{i}$)
  • $\mathsf{H}_{4}$: Simulate responses from $\tilde{\mathbf{y}}_{i}$ and simulated key shares
@@
Lemma P.8 (Pre-signing Simulation (ROM)).
$\Pr[\mathsf{H}_{2}=1]=\Pr[\mathsf{H}_{3}=1]$ (perfect indistinguishability in the random oracle model)
Proof P.9.
We work in the UC random oracle model: $H$ is modeled as an ideal functionality $\mathcal{F}_{\mathsf{RO}}$ (following [Canetti01, BR93]). The simulator $\mathcal{S}^{\mathsf{P3+}}$ programs $\mathcal{F}_{\mathsf{RO}}$ at honest parties’ pre-image inputs: when the simulator generates $\tilde{\mathsf{Com}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$, it programs $H(\tilde{\mathbf{y}}_{i}\|\mathbf{A}\tilde{\mathbf{y}}_{i}\|\mathsf{set\_id}):=\tilde{\mathsf{Com}}_{i}$ at the simulated pre-image (the adversary does not know $\tilde{\mathbf{y}}_{i}$ and cannot query $H$ on this input without knowing the simulation randomness).
Pre-signing values visible to the adversary:
  • $\mathsf{Com}_{i}=H(\mathbf{y}_{i}\|\mathbf{w}_{i}\|\mathsf{set\_id})$: In the ROM, $\mathsf{Com}_{i}$ is uniform over $\{0,1\}^{256}$ regardless of $\mathbf{y}_{i}$’s distribution, since the adversary does not know $\mathbf{y}_{i}$ (or $\mathbf{w}_{i}=\mathbf{A}\mathbf{y}_{i}$) and cannot query $\mathcal{F}_{\mathsf{RO}}$ at the pre-image without knowing the honest party’s nonce. (Note: in the real protocol with $T\geq 2$, $\mathbf{y}_{i}=\mathbf{F}(i)$ is $\mathbb{Z}_{q}$-uniform: the high-degree DKG coefficients (degree $\geq 1$) are sampled $R_{q}$-uniformly, and for evaluation point $i\neq 0$ their contribution is uniform by additive mixing, giving $\mathbf{y}_{i}\sim\mathbb{Z}_{q}$-uniform exactly. When $T=1$ there are no high-degree terms and $\mathbf{y}_{i}=\mathbf{y}_{0}$ is the constant nonce, but this is a degenerate threshold that is excluded by the protocol requirement $T\geq 2$. The Irwin-Hall distribution applies to the aggregate nonce $\mathbf{y}=\mathbf{F}(0)=\sum_{k}\hat{\mathbf{y}}_{k}$, not to individual shares $\mathbf{y}_{i}=\mathbf{F}(i)$.) The simulated $\tilde{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}$ and $\tilde{\mathsf{Com}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ therefore match the real distributions exactly (SD =0=0).
  • $\mathbf{W}_{i}=\lambda_{i}\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)}$ (masked commitment sent to CP combiner): In P3+, $\mathbf{W}_{i}$ is transmitted only to the CP combiner, not broadcast to other parties. By the 1-of-2 CP honest assumption, at least one CP is honest; in this proof we assume the adversary does not control the combiner’s input-aggregation role (if a CP is corrupted, the relevant 2PC simulation case handles it below). Since $\mathbf{W}_{i}$ never reaches the adversary, no mask-hiding argument is required. The value is hidden entirely by the honest combiner. (Note: Lemma 4.15 applies to the P2 broadcast model where $\mathbf{W}_{i}$ is sent to all parties; it does not apply here.)
Thus:
@@
Proof P.11.
The simulator computes each honest party’s response deterministically from its simulated DKG values, producing an identical distribution to the real protocol:
@@
where $\mathbf{y}_{i}$ are the nonce shares from the simulated nonce DKG (Lemma P.8) and $\mathbf{s}_{1,i},\mathbf{s}_{2,i}$ are the key shares from the simulated key DKG. Since the simulator possesses the actual simulated values (not just their distributions), it computes $\mathbf{z}_{i}$ and $\mathbf{V}_{i}$ using the same deterministic function as the real protocol; the resulting joint distribution is identical to the real execution. For $\mathbf{V}_{i}$, the simulator computes the masked r0-check share using the simulated key shares and the PRF masks (derived from the same seeds used in earlier hybrids). Perfect indistinguishability follows: $\Pr[\mathsf{H}_{3}=1]=\Pr[\mathsf{H}_{4}=1]$.
@@
Protocol $\Pi_{r_{0}}^{\mathsf{P3+}}$ UC-realizes $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ in the $(\mathcal{F}_{\mathsf{OT}},\mathcal{F}_{\mathsf{GC}})$-hybrid model against static malicious adversaries corrupting up to T1T-1 signers and at most one of $\{\mathsf{CP}_{1},\mathsf{CP}_{2}\}$, with security-with-abort. Key DKG is handled by a trusted dealer (modeled as a trusted setup phase); the simulator acts as the trusted dealer and directly generates all key shares.
@@
$|\Pr[\mathsf{REAL}=1]-\Pr[\mathsf{IDEAL}=1]|\leq\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{ext}}<2^{-\kappa}$
where $\kappa$ is the computational security parameter, $\epsilon_{\mathsf{OT}}$ is the OT error, $\epsilon_{\mathsf{GC}}$ is the garbled-circuit simulation error (over honest garbler [BHR12]), and $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$ is the cut-and-choose extraction error when $\mathsf{CP}_{1}$ is corrupted (LP07 extractability [LP07] with $B$ circuits, $B/2$ opened for verification).
Proof P.13.
@@
$|\Pr[\mathsf{H}_{0}]-\Pr[\mathsf{H}_{5}]|$ $\leq\sum_{i=0}^{4}|\Pr[\mathsf{H}_{i}]-\Pr[\mathsf{H}_{i+1}]|+|\Pr[\mathsf{H}_{2.5,\mathsf{ext}}]-\Pr[\mathsf{H}_{3}]|$
$=\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{ext}}+0+0+0$